# CaDRE: Controllable and Diverse Generation of Safety-Critical Driving Scenarios using Real-World Trajectories

### v1

@@ III-C Occupancy-Aware Restart
As $T\rightarrow+\infty$, OAR degenerates to the uniform sampling. With a lower $T$, OAR assigns a higher probability to those elites who have more empty neighbors. For efficient implementation, we use a 3D convolution kernel to compute the number of empty cells around each elite.
Figure 2: Illustration of Occupancy-Aware Restart.
@@
TABLE I: The final performance of coverage, mean objective, QD score. We report the mean and variance over 5 perturbed vehicles for each scene. The QD score is shown in multiples of $31e$ 3.
@@
Unprotected cross-turn High-speed lane-change U-turn Method Coverage ($\uparrow$) Mean Obj ($\uparrow$) QD Score ($\uparrow$) Coverage ($\uparrow$) Mean Obj ($\uparrow$) QD Score ($\uparrow$) Coverage ($\uparrow$) Mean Obj ($\uparrow$) QD Score ($\uparrow$) Random 0.140$\pm$0.021 0.499$\pm$0.123 0.285$\pm$0.098 0.310$\pm$0.158 0.310$\pm$0.158 0.209$\pm$0.131 0.188$\pm$0.066 0.381$\pm$0.134 0.320$\pm$0.151 CMA-ES 0.182$\pm$0.033 0.672$\pm$0.076 0.489$\pm$0.090 0.163$\pm$0.018 0.540$\pm$0.143 0.347$\pm$0.086 0.228$\pm$0.118 0.447$\pm$0.228 0.502$\pm$0.253 REINFORCE 0.210$\pm$0.031 0.649$\pm$0.115 0.551$\pm$0.155 0.488$\pm$0.144 0.488$\pm$0.144 0.472$\pm$0.161 0.286$\pm$0.010 0.641$\pm$0.117 0.731$\pm$0.117 SVPG 0.219$\pm$0.032 0.607$\pm$0.130 0.553$\pm$0.155 0.438$\pm$0.096 0.438$\pm$0.096 0.437$\pm$0.166 0.290$\pm$0.020 0.577$\pm$0.120 0.665$\pm$0.110 CaDRE (ours) 0.565$\pm$0.054 0.829$\pm$0.062 1.884$\pm$0.309 0.541$\pm$0.079 0.627$\pm$0.142 1.375$\pm$0.436 0.542$\pm$0.094 0.793$\pm$0.073 1.699$\pm$0.219
@@
Figure 3: Coverage and QD score v.s. the number of samples. The solid lines represent the mean, and the shaded area presents the standard deviation over 5 perturbed vehicles.
@@ IV-A Experimental Setup
Real-world Trajectories. We pick three representative scenarios from nuPlan v1.1 [50]: unprotected cross-turn, high-speed lane-change, and U-turn. All scenarios have a time horizon of $150$ frames @ $10Hz$ and are down-sampled to 5Hz in our experiments.
Reactive Ego Policy. We implement a rule-based ego policy: The ego vehicle will follow the reference trajectory. However, if there is a vehicle within $5m$ and $[-\pi/4,\pi/4]$ of the body frame of the ego, the ego vehicle will brake at $-7m/s^{2}$ and maximum steering angle ±$\pm\pi/8$ depending on the relative position of the vehicle w.r.t. the body frame of the ego. Recall that the problem is formulated as a black-box optimization, and CaDRE is agnostic to the ego agent’s policy.
@@
Evaluation Metrics. We focus on three criteria that measure the quality and diversity of the archive, which are standard metrics in the QD literature [49, 39, 51, 52].
  • Coverage $\in[0,1]$: Proportion of cells in the archive that have an elite.
@@
Baselines. We study a mixture of sampling- and RL-based methods that have been employed by the existing literature.
@@
  • CMA-ES [53]: CMA-ES iteratively updates a population of solutions based on their fitness, using statistical information from previous generations to adaptively adjust the search distribution towards optimal regions of the solution space, on which CMA-ME is based but without QD.
@@
  • Stein Variational Policy Gradient (SVPG) [55]: SVPG is an improved version of multi-particle REINFORCE. SVPG introduces a maximum entropy policy optimization framework that explicitly encourages diverse solutions and better exploration. Similar to multi-particle REINFORCE, we set the number of particles to be the same as the batch size of CMA-ME.
We use the QD algorithm library pyribs [39] to implement our framework. We aim to answer the following questions in our experimental study:
@@ IV-B Sample-efficiency Compared with Baseline Methods
The coverage and QD score v.s. samples are shown in Figure 3. CaDRE outperforms all baselines with significant margins in three different scenarios, which demonstrates that CaDRE discovers not only high-quality but also diverse scenarios much faster than Random Search, SVPG, and REINFORCE, with the same number of samples. CaDRE utilizes the Covariance Matrix Adaptation (CMA) strategy, which adapts the search distribution over generations to increase the likelihood of sampling promising areas of the solution space. This adaptation is based on information from previous generations, allowing CaDRE to focus its sampling on regions with higher potential for high-quality solutions. Unlike Random Search, which samples uniformly across the solution space without learning from previous samples, CaDRE dynamically narrows its search to more promising regions. CMA-ES, despite being on which CMA-ME is based, shows a completely different purpose, which is to the likelihood of increasing objective and, therefore, quickly converges to a single optimum. SVPG, and REINFORCE, while more directed than Random Search, may still struggle with efficiently exploring complex problem spaces due to their focus on gradient-based optimization.
Table I shows the final performance of coverage, mean objective, and QD score. With the same number of samples in unprotected cross-turn, CaDRE achieves $158.0\%$ more coverage, $36.6\%$ higher mean objective, leading to a $240.7\%$ improvement in QD score than the best-performing baseline SVPG. It again highlights the superior exploration and exploitation capability of CaDRE compared to the baselines. Table II shows the ablation of OAR in the high-speed lane-change scenario. OAR improves the QD score of individual vehicles by a maximum of $33.0\%$, which demonstrates the effectiveness of OAR.
TABLE II: QD score of occupancy-aware restart with different temperatures. The QD score is shown in multiples of $31e$ 3. We include the percentage improvement w.r.t $01/T=$ 0 in parentheses.
Index $01/T=$ 0 $51/T=$ 5 $101/T=$ 10 1 1.808 1.855 (2.6%) 2.026 (12.0%) 2 0.519 0.639 (22.9%) 0.691 (33.0%) 3 1.277 1.133 (-11.3%) 1.229 (-3.7%) 4 1.393 1.515 (8.8%) 1.367 (-1.9%) 5 1.444 1.299 (-10.0%) 1.563 (8.2%)
@@ IV-C Analysis of the Generated Safety-Critical Scenarios
Visualization of Archives. We visualize the final archives in Fig. 6. It is observed that the proposed CaDRE leads to a much higher occupancy as well as mean objective in the final archives than the baseline SVPG. The main reason is that CaDRE explicitly encourages sustained exploration throughout optimization. CaDRE employs CMA-ME, which is particularly adept at exploring complex landscapes and finding a large amount of high-quality scenarios in different measures. Although the repulsive force in SVPG indeed introduces diversity among the particles to avoid premature convergence to local optima, the primary focus remains on optimizing a solution rather than explicitly seeking out diverse solutions across a range of measures.
@@
Distribution of Measure Values. Figure 4 visualizes the distribution of the measure function values. Compared to SVPG, CaDRE generates a denser and wider range of measure function values. However, both methods struggle to find safety-critical scenarios with very little steering perturbation, which is reasonable as the original scenarios only contain safe and regular traffic, and the perturbation is bounded.
Visualization of Generated Scenarios. In Fig. 5, we visualize five generated scenarios for the unprotected cross-turn, high-speed lane-change, and U-turn, respectively. The visualization shows that we are able to retrieve diverse critical scenarios in a controllable manner. For example, in the unprotected cross-turn, we can control the perturbed vehicle hitting the right side of the ego vehicle by steering a little bit from the original trajectory or hitting the left side by overtaking from the left, simply by asking for different combinations of measure function values $[m_{1},m_{2},m_{3}]$ in the archive.
IV-D Limitations and Future Directions
Although CaDRE has demonstrated the ability to generate diverse and controllable scenarios with superior sample efficiency, it is not without its limitations. First, CaDRE does not consider the lane information and road conditions such as barriers. It could generate scenarios that are kinematically feasible but unlikely in real life, such as going through the median of the highway. Second, CaDRE only perturbs one of the background vehicles. However, in the real world, there exist some critical scenarios induced by more than one vehicle or not directly caused by a collision with the perturbed vehicle. For example, a background vehicle makes a lane change to avoid hitting another vehicle breaking in front, thus hitting the ego vehicle from the side. It is a promising direction to extend CaDRE to consider road information and perturb more than one vehicle or one vehicle that indirectly causes the collision.
V CONCLUSIONS
In this work, we develop a framework CaDRE for generating safety-critical scenarios. As a variant of the QD algorithm, CaDRE enhances the diversity and controllability of the scenario generation process, thus providing an effective instrument for the simulation-based assessment of autonomous vehicles. We conduct extensive experiments on three representative scenarios: unprotected cross-turn, high-speed lane-change, and U-turn. The experimental results show that CaDRE can generate and retrieve diverse and high-quality scenarios with better sample efficiency compared with both RL- and sampling-based methods.
ACKNOWLEDGMENT
This work was partially performed during PH’s internship at the Bosch Center for Artificial Intelligence; we thank Bosch for the use of computing resources. The authors additionally thank Uksang Yoo and Benjamin Stoler for valuable conversations. The authors gratefully acknowledge the support from the National Science Foundation under grants CNS-2047454.

### v2

@@ III-C CaDRE Algorithm
Illustrated in Fig. 1(b), OAR considers the occupancy rate of neighboring cells and assigns a higher probability to elites with more empty neighboring cells. More specifically, given the neighbor empty rate of $L$ elites $r_{1},\ldots,r_{L}$ and the temperature $\tau$, the softmax probability of restarting from elite $i$ is computed by $p_{i}=\frac{e^{r_{i}/\tau}}{\sum_{j=1}^{L}e^{r_{j}/\tau}}$. OAR degenerates to the uniform sampling as $\tau\to\infty$. With a lower $\tau$, OAR assigns a higher probability to those elites who have more empty neighbors. For an efficient implementation, we use a 3D convolution kernel to compute the number of empty cells around each elite.
@@
TABLE I: The final performance of coverage, mean objective, QD score. We report the mean and variance over 5 perturbed vehicles for each scene. The QD score is shown in multiples of $1000$.
@@
Unprotected cross-turn High-speed lane-change U-turn Method Coverage ($\uparrow$) Mean Obj ($\uparrow$) QD Score ($\uparrow$) Coverage ($\uparrow$) Mean Obj ($\uparrow$) QD Score ($\uparrow$) Coverage ($\uparrow$) Mean Obj ($\uparrow$) QD Score ($\uparrow$) Random 0.140$\pm$0.021 0.499$\pm$0.123 0.285$\pm$0.098 0.162$\pm$0.024 0.310$\pm$0.158 0.209$\pm$0.131 0.188$\pm$0.066 0.381$\pm$0.134 0.320$\pm$0.151 GOOSE 0.279$\pm$0.097 0.368$\pm$0.058 0.429$\pm$0.180 0.257$\pm$0.082 0.385$\pm$0.158 0.443$\pm$0.291 0.141$\pm$0.082 0.490$\pm$0.158 0.325$\pm$0.252 CMA-ES 0.182$\pm$0.033 0.672$\pm$0.076 0.489$\pm$0.090 0.163$\pm$0.018 0.540$\pm$0.143 0.347$\pm$0.086 0.228$\pm$0.118 0.447$\pm$0.228 0.502$\pm$0.253 REINFORCE 0.210$\pm$0.031 0.649$\pm$0.115 0.551$\pm$0.155 0.243$\pm$0.033 0.488$\pm$0.144 0.472$\pm$0.161 0.286$\pm$0.010 0.641$\pm$0.117 0.731$\pm$0.117 SVPG 0.219$\pm$0.032 0.607$\pm$0.130 0.553$\pm$0.155 0.265$\pm$0.026 0.410$\pm$0.129 0.437$\pm$0.166 0.290$\pm$0.020 0.577$\pm$0.120 0.665$\pm$0.110 CaDRE (ours) 0.565$\pm$0.054 0.829$\pm$0.062 1.884$\pm$0.309 0.541$\pm$0.079 0.627$\pm$0.142 1.375$\pm$0.436 0.542$\pm$0.094 0.793$\pm$0.073 1.699$\pm$0.219
@@
Figure 2: Coverage and QD score v.s. number of samples. Solid lines represent the mean, and the shaded area presents the standard deviation over 5 perturbed vehicles.
@@ IV-A Experimental Setup
Real-world Trajectories. We pick three representative scenarios from nuPlan v1.1 [49]: unprotected cross-turn, high-speed lane-change, and U-turn. All scenarios have a total time horizon of 15 seconds.
Reactive Ego Policy. Following [12, 31], we assume that the ego vehicle is reactive to nearby vehicles. We implement a rule-based ego policy: The ego vehicle will try to follow the reference trajectory, and if there is a vehicle within $5m$ and $[-\pi/4,\pi/4]$ of the body frame of the ego vehicle, it will brake at $-7m/s^{2}$ and steer away from that vehicle with maximum steering angle ±$\pm\pi/8$. Since the problem is formulated as a black-box optimization, CaDRE does not require access to the reactive policy and would work with any other ego policies.
@@
Evaluation Metrics. We focus on three criteria that measure the quality and diversity of the generated scenarios, which are standard metrics in the QD literature [38, 19, 39, 50].
  • Coverage $\in[0,1]$: Proportion of cells in the archive that have an elite. This is a measure of diversity of the generated scenarios. Note that while coverage is a standard metric for QD, there lacks such a principled metric to evaluate diversity in safety-critical scenario generation literature.
@@
Baselines. We study a mixture of sampling- and RL-based methods that have been employed by existing literature.
@@
  • GOOSE [51]: a state-of-the-art RL-based approach for safety-critical scenario generation. It parameterizes adversarial trajectories as non-uniform rational B-splines (NURBS) and iteratively modifies their control points for diverse scenario generation. Note that this is also an example of diversity by construction, where diversity is promoted through the representation of scenarios.
  • CMA-ES [52]: CMA-ES is an evolutionary strategy that iteratively updates a population of solutions based on their fitness and adaptively adjusts the search distribution towards optimal regions of the solution space.
@@
  • Stein Variational Policy Gradient (SVPG) [54]: SVPG is an improved version of multi-particle REINFORCE by using a maximum entropy policy optimization framework that explicitly encourages diverse solutions and better exploration. The number of particles is the same as the batch size of CaDRE.
We adopt the QD algorithm library pyribs [19] to implement our framework. We aim to answer the following questions in our experimental study:
@@ IV-B Sample-efficiency Compared with Baseline Methods
The coverage and QD score v.s. samples are shown in Figure  2. CaDRE outperforms all baselines with significant margins in three different scenarios, which demonstrates that CaDRE discovers both high-quality and diverse scenarios more sample-efficiently than Random Search, GOOSE, SVPG, and REINFORCE. Under the hood, CaDRE adapts the search distribution over generations to increase the likelihood of sampling in promising areas of the solution space, in contrast to Random Search, which samples uniformly across the solution space. CMA-ES maximizes the likelihood of increasing the objective and thus quickly converges to a single optimum. SVPG and REINFORCE, while exploring more than CMA-ES, still struggle with exploring complex problem spaces due to their focus on gradient-based optimization. Finally, GOOSE, a dedicated method for safety-critical scenario generation, claimed that the NURBS parameterization of traffic scenarios is sufficient for diversity. However, it is clear from the results that its diversity, as measured by coverage, lags far behind CaDRE, though being competitive with other baselines.
@@
Table I shows the final performance of coverage, mean objective, and QD score. With the same number of samples, CaDRE achieves $158.0\%,122.6\%,89.5\%$ more coverage, $36.6\%,28.5\%,23.7\%$ higher mean objective, leading to a $240.7\%,191.3\%,132.4\%$ improvement in QD score than the best-performing baselines in three representative scenarios, respectively. It again highlights the superior exploration and exploitation capability of CaDRE compared to the baselines.
Table II shows the ablation of OAR in the high-speed lane-change scenario. $\tau\to\infty$ is equivalent to QD algorithms without OAR. OAR improves the QD score of individual vehicles by a maximum of $33.0\%$, which demonstrates the effectiveness of OAR. We also observe that the biggest improvements come from harder problems (low QD score). The intuition is that when it is harder to find a safety-critical scenario, then a more effective exploration strategy provides a more significant performance boost.
TABLE II: QD score of occupancy-aware restart with different temperatures. The QD score is shown in multiples of $1000$. The percentage improvements w.r.t $\tau\to\infty$ are in parentheses.
Index $1$ $2$ $3$ $4$ $5$ $\tau\to\infty$ 1.808 0.519 1.277 1.393 1.444 $\tau=1/5$ 1.855 (2.6%) 0.639 (22.9%) 1.133 (-11.3%) 1.515 (8.8%) 1.299 (-10.0%) $\tau=1/10$ 2.026 (12.0%) 0.691 (33.0%) 1.229 (-3.7%) 1.367 (-1.9%) 1.563 (8.2%)
@@ IV-C Analysis of the Generated Safety-Critical Scenarios
Visualization of Archives. We visualize the final archives in Fig. 5. It is observed that the proposed CaDRE leads to a much higher occupancy as well as mean objective in the final archives than the best-performing baseline SVPG. The main reason is that CaDRE explicitly encourages sustained exploration throughout optimization. Although the repulsive force in SVPG indeed introduces diversity among the particles to avoid premature convergence to local optima, the primary focus remains on optimizing a solution rather than explicitly seeking out diverse solutions across a range of measures.
@@
Visualization of Generated Scenarios. In Fig. 4, we visualize five generated scenarios for the unprotected cross-turn, high-speed lane-change, and U-turn, respectively. The visualization shows that we are able to generate and retrieve diverse critical scenarios in a controllable manner. For example, in the unprotected cross-turn, we can control the perturbed vehicle hitting the right side of the ego vehicle by steering a little bit from the original trajectory or hitting the left side by overtaking from the left, simply by asking for different combinations of measure function values $[m_{1},m_{2},m_{3}]$ in the archive.
V CONCLUSIONS AND DISCUSSION
In this work, we develop a framework CaDRE for generating safety-critical scenarios. CaDRE enhances the diversity and controllability of the scenario generation process while retaining realism. We conduct extensive experiments on three representative scenarios: unprotected cross-turn, high-speed lane-change, and U-turn. The experimental results show that CaDRE can generate and retrieve diverse and high-quality scenarios with better sample efficiency compared to existing methods.
We envision many extensions possible to CaDRE. At the moment, CaDRE only perturbs one background vehicles at a time. Alternative parameterization of traffic scenarios or objective and measure functions can make CaDRE more general or suited to specific settings. Second, CaDRE does not consider the lane information and road conditions such as barriers. It is possible to incorporate road information to CaDRE by augmenting the existing simulator or adding a post-hoc pruning step based on the drivable area.